Johns mother is 5 times as old as john is. Three years ago She was 9 times his age. Use simultaneous equations to solve

Answer by macston(5194)

J=John's age; M=mother's age=5J
M-3=9(J-3) Substitute for M.
5J-3=9J-27 Subtract 5J from each side.
-3=4J-27 Add 27 to each side.
24=4J Divide each side by 4.
6=J ANSWER 1: John is 6 years old.
M=5J=5(6yrs)=30yrs ANSWER 2: Mother is 30 years old.
.
CHECK:
.
M-3=9(J-3)
30-3=9(6-3)
27=9(3)
27=27